### v1

@@ Abstract
Cell imaging assays utilizing fluorescence stains are essential for observing sub-cellular organelles and their responses to perturbations. Immunofluorescent staining process is routinely in labs, however the recent innovations in generative AI is challenging the idea of IF staining are required. This is espeically true when the availability and cost of specific fluorescence dyes is a problem to some labs. Furthermore, staining process takes time and leads to inter-/intra-technician and hinders downstream image and data analysis, and the reusablity of image data for other projects. Recent studies showed the use of generated synthetic immunofluorescence (IF) images from brightfield (BF) images using generative AI algorithms in the literature. Therefore in this study, we benchmark and compare five models from three types of IF generation backbones—CNN, GAN, and diffusion models—using a publicly available dataset. This paper not only serves as a comparative study to determine the best-performing model but also proposes a comprehensive analysis pipeline for evaluating the efficacy of generators in IF image synthesis. We highlighted the potential of deep learning-based generators for IF image synthesis, while also discussed potential issues and future research directions. Although generative AI shows promise in simplifying cell phenotyping using only BF images with IF staining, further research and validations are needed to address the key challenges of model generalisability, batch effects, feature relevance and computational costs.
@@ 1 Introduction
High-throughput screening (HTS) has emerged as a critical method in drug discovery, enabling a fast evaluation of vast libraries of compounds for their therapeutic potential. In this context, cell image assays that utilize fluorescence microscopy are of great importance, as they allow scientists to observe and analyse cellular responses to pharmaceutical compounds. By applying fluorescent markers to highlight specific cellular structures, these assays yield detailed information about the effects of drugs on cell shape and function.
However, HTS via cell imaging assays presents distinct challenges, particularly in data acquisition and analysis.The first issue is the resource efficiency. Fluorescence microscopy requires specific dyes or labels, which can be expensive or limited in availability. In labs, it happens when certain fluorescent stains are not available, which inconveniently delays experiments. Additionally, the cost of fluorescence painting is also a concern. Specific dyes are very expensive, hindering scientific research especially for smaller labs, such as the nucleic acid dye TOTO-3 iodide, which has strong binding affinity for dsDNA. Finally, some specific fluorescent dyes introduce cytotoxicity to live cells and alter endogenous cell function and phenotype. Synthesizing fluorescence images negates this issue, allowing for non-destructive analysis of cells and detection of subtle organelle signals. Brightfield imaging, although easier to acquire, uses light and stains to visualize general tissue morphology and cell structure lacking of detailed cellular and molecular analysis
@@
We aim to establish a foundation for the broader application of generative AI using cell imaging assays in preclinical settings. Our work showcases the potential of synthetic data in improving the efficiency of cell phenotyping analysis using only brightfield images. By addressing current research gaps, our project strives to set a new standard for using AI to predict IF staining. Ultimately, this work could significantly speedup and streamline workflows in biomedical research. Finally we conclude this research by answering the question that GenAI cannot replace IF stain.
@@ 2 Related Work in the Synthesis of Pseudo-IF Images
More recently, diffusion models [11] have emerged as a novel technology in image synthesis. These models add noise to images iteratively and then reverse this process to restore the original images with high fidelity. Diffusion models outperform GANs [7] in terms of fidelity and diversity due to their ability to preserve intricate details. Their effectiveness specifically in synthesizing IF images was validated in [5].
@@ 3.1 Dataset
Our model was trained on the public cpg0000 dataset [3] from the Cell Painting Gallery  [^1] https://registry.opendata.aws/cellpainting-gallery/, which consists of over three million pairs of BF and IF images. This dataset includes samples treated with over 300 compounds, 160 CRISPR gene knockouts in both A549 and U2OS cell backgrounds at two time points. To evaluate the generalisability of our model, we selected a subset of these images, focusing on eight plates exposed to various compounds, as detailed in Table 1. Each plate contains 384 wells, with each well either treated with a specific compound or DMSO as a control. Two plates were used for training, one for validation and one for testing. Additionally, two plates were added for evaluating the generalisability of the generative AI models in response to different type of perturbations.
@@
The size of each square image is 645 $\mu m$ (1080 pixels in both height and width). During training and inference, we resized all images to $512\times 512$ pixels.
@@ 3.2 Synthesis Models Compared in This Study
3.2.1 CNN Model: UNet Architecture [17]
For the CNN model tasked with synthesizing IF images from BF images, we have chosen the UNet architecture. The UNet model is characterized by its encoder-decoder structure: the encoder compresses the input image into a feature-rich, compact representation while the decoder then progressively reconstructs the target image from the encoded representation. UNets are particularly noted for their skip connections, which bypass the encoding layers, feeding information directly to the corresponding decoding layers. These connections are essential as they carry fine-grained spatial information across the network, aiding in the precise localisation needed for accurate synthesis. Mean Squared Error (MSE) loss was utilised to train the UNet, which measures the average squared difference between the estimated values and what is estimated.
@@ 3.2.2 GAN Models
We have implemented two variations of GANs differentiated by how they incorporate the input BF image condition:
@@ 3.2.3 Diffusion Models
Palette [18]. Within the diffusion model framework, Palette utilises a U-Net architecture for the denoising process; the input BF images are concatenated with noisy images, providing conditional context throughout the iterative denoising steps. As with the CNN UNet implementation, MSE loss is used to guide the denoising process towards producing high-quality IF images.
@@ 3.3 Evaluation Methods
Various algorithms have been proposed in previous studies to evaluate synthetic images [10, 19, 1]. In this paper, we define two categories of metrics: image quality evaluation metrics and biological meaning evaluation metrics.
@@ 3.3.2 Biological Meaning Metrics: Feature Correlation and MoA Prediction Accuracy
The first evaluation metric is the correlation between these features. The mean values for each feature group were organized into correlation matrices and visualized as heatmaps. In CellProfiler, features from each object were divided into several categories: area/shape, colocalization, granularity, intensity, neighbours, radial distribution, and texture.
The second evaluation metric is the predictive performance of these single-cell features for determining the mechanism of action (MoA) of perturbations e.g. compounds. The MoA refers to the biological process through which a compound achieves its pharmacological effect, such as targeting specific proteins or modulating certain pathways. Understanding a compound’s MoA is critical in pre-clinical phases as it helps identify potential efficacy and toxicity effects.
@@ 4.1 Pix2Pix Model Performs the Best Regarding Image Quality Evaluation
Visual Comparison. Through sufficient training, all five models captured the distinctive style of the IF images, though, there are noticeable differences in details between the generated images and the ground truth images. Among these models, we discovered that the Pix2Pix model captures the details of the real IF images the most accurately.
@@
In Figure 3(a), when comparing the MSE values across different channels, we noticed that, with the exception of SPADE, the MSEs for the DNA channel in all models are significantly lower ($p<0.05$) than those for other channels. This suggested that predicting the cell nucleus channel is easier to achieve than transferring the styles of BF or searching for subtle and embeded subcellular structures in other channels (Mito, AGP, ER, and RNA). This is likely because nuclei have a clearer shape and more defined semantics compared to other structures. To improve synthesis across all structures, additional conditional mechanisms may be required during the synthesis process across all model families.
@@ 5.1 Three Typical Artifacts in Synthetic IF Images
In Figure 2, we highlighted three types of artifacts that were identified when generative AI predicted IF images from BF images: 1) Blurring in CNN generated images (Figure 2c): This artifact is prevalent in the CNN-predicted IF images through all channels. This is possibly attributed to the model’s MSE-based optimization loss function. The MSE loss encourages minimising pixel-wise differences, blurring fine details and leading to overly smoothed images. 2) Collapsed Inner Structure especially in the AGP channel: Some models fail to accurately replicate the structure, creating hallucinations or "collapsed biological structures." This hallucination is likely due to the model’s inability to condition the synthesis procedure upon given BF conditions. 3) Intensity Shift and High Background Noises: In diffusion models, the predicted images often have significant background noise and intensity shifts. These artifacts can occur due to the nature of the diffusion process, where each step requires intensity clipping that may introduce noise and modify image brightness.
@@ 5.3 Problematic MSE
where $y_{i}$ is the ground truth pixel value and $\hat{y}_{i}$ is the predicted pixel value.
However, MSE loss is limited as an evaluation metric as it only considers the pixel value differences and aggregates the whole image regardless of background and signals. In a biological context, this may pose serious potential issues in downstream applications. In Figure 7, we present two synthetic IF images (2,3) alongside their corresponding ground truth image (1). Figure 7 (2a) was generated by the Pix2Pix network while Figure 7 (3a) was produced by intensity reduction. As shown in Figure 7 (2b), while Figure 7 (2a) has a smaller image pixel intensity difference compare to the real image, and therefore a lower MSE loss, the cell morphology (in terms of nuclear size, shape and texture) is incorrectly predicted, leading to an inaccurate estimation of cell mortality, shown with two red arrows. On the other hand, Figure 7 (3b) largely retains nuclear morphology despite large differences in stain intensity compared to the real image. Thus, while MSE, PSNR, and SSIM can roughly assess image intensity, they do not fully capture the biological relevance. As such, leveraging metrics that assess the biological significance of these synthetic images is crucial for robust model evaluation.
@@ 5.4 Discrepancy in Feature Correlation and MoA Prediction Accuracy
As illustrated, the single-cell features identified in CNN and GAN-based models closely follow the real feature manifold, which is consistent with the image quality results. In contrast, the features of cells generated by diffusion models occupy a different space and exhibit different characteristics. However, this does not imply that these out-of-real data points lack classification ability. Indeed, we observed a significant location difference between the DMSO group and the inhibitor-treated group in features generated by the diffusion models, as highlighted by the red and purple points in Figure 8.
@@
5.5 Lack of Generalisability to Unseen Cell Lines

### v2

@@ Abstract
Cell imaging assays utilizing fluorescence stains are essential for observing sub-cellular organelles and their responses to perturbations. Immunofluorescent staining process is routinely in labs, however the recent innovations in generative AI is challenging the idea of wet lab immunofluorescence (IF) staining. This is especially true when the availability and cost of specific fluorescence dyes is a problem to some labs. Furthermore, staining process takes time and leads to inter-intra-technician and hinders downstream image and data analysis, and the reusability of image data for other projects. Recent studies showed the use of generated synthetic IF images from brightfield (BF) images using generative AI algorithms in the literature. Therefore, in this study, we benchmark and compare five models from three types of IF generation backbones—CNN, GAN, and diffusion models—using a publicly available dataset. This paper not only serves as a comparative study to determine the best-performing model but also proposes a comprehensive analysis pipeline for evaluating the efficacy of generators in IF image synthesis. We highlighted the potential of deep learning-based generators for IF image synthesis, while also discussed potential issues and future research directions. Although generative AI shows promise in simplifying cell phenotyping using only BF images with IF staining, further research and validations are needed to address the key challenges of model generalisability, batch effects, feature relevance and computational costs.
@@ 1 Introduction
High-throughput screening (HTS) has emerged as a critical method in drug discovery, enabling the fast evaluation of vast libraries of compounds for their therapeutic potential. In this context, cell image assays that utilise fluorescence microscopy are of great importance, as they allow scientists to observe and analyse cellular responses to pharmaceutical compounds. By applying fluorescent markers to highlight specific cellular structures, these assays yield detailed information about the effects of drugs on cell shape and function.
However, HTS via cell imaging assays presents distinct challenges, particularly in data acquisition and analysis. The first issue is the resource efficiency. Fluorescence microscopy requires specific dyes or labels, which can be expensive or limited in availability. In labs, certain fluorescent stains are often not immediately available, which inconveniently delays experiments. Additionally, the cost of fluorescence painting is also a concern. Specific dyes, such as the nucleic acid dye TOTO-3 iodide that has strong binding affinity for dsDNA, are very expensive, hindering research feasibility and scale especially for smaller labs. Finally, some specific fluorescent dyes introduce cytotoxicity to live cells and alter endogenous cell function and phenotype. Synthesising fluorescent images negates this issue, allowing for non-destructive analysis of cells and detection of subtle organelle signals. Brightfield (BF) imaging uses light and stains to visualise general tissue morphology and cell structure. Although easier to acquire, the images are insufficient for detailed cellular and molecular analysis
@@
We aim to establish a foundation for the broader application of generative AI using cell imaging assays in preclinical settings. Our work showcases the potential of synthetic data in improving the efficiency of cell phenotypic analysis using only brightfield images. By addressing current research gaps, our project strives to set a new standard for using AI to predict IF staining. Ultimately, this work could significantly accelerate and streamline workflows in biomedical research. Finally, we conclude this research by answering the question that GenAI cannot entirely replace IF staining at present.
@@ 2 Related Work in the Synthesis of Pseudo-IF Images
More recently, diffusion models [12] have emerged as a novel technology in image synthesis. These models iteratively add noise to images and then reverse this process to restore the original images with high fidelity. Diffusion models outperform GANs [8] in terms of fidelity and diversity due to their ability to preserve intricate details. Their effectiveness specifically in synthesizing IF images was validated in [6].
@@ 3.1 Dataset
Our model was trained on the public cpg0000 dataset [4] from the Cell Painting Gallery  [^1] https://registry.opendata.aws/cellpainting-gallery/, which consists of over three million pairs of BF and IF images. This dataset includes samples treated with over 300 compounds, 160 CRISPR gene knockouts in both A549 and U2OS cell backgrounds at two time points. To evaluate the generalisability of our model, we selected a subset of these images, focusing on eight plates exposed to various compounds, as detailed in Table 1. Each plate contains 384 wells, with each well either treated with a specific compound or DMSO as a control. Two plates were used for training, one for validation and one for testing. Additionally, two plates under different pertubations were added for evaluating the generalisability of generative models.
@@
The size of each square image is 645 $\mu m$ (1080 pixels in both height and width). During training and inference, we resized all images to $512\times 512$ pixels. We enhanced the contrast by cutting off the peak and bottom 5% of pixels in the histogram. All images were saved in the 0-255 range during evaluation and normalized to the 0-1 range during training.
@@ 3.2 Synthesis Models Compared in This Study
3.2.1 CNN Model
For the CNN model tasked with synthesizing IF images from BF images, we have chosen the UNet architecture [20]. The UNet model is characterized by its encoder-decoder structure: the encoder compresses the input image into a feature-rich, compact representation while the decoder then progressively reconstructs the target image from the encoded representation. UNets are particularly noted for their skip connections, which bypass the encoding layers, feeding information directly to the corresponding decoding layers. These connections are essential as they carry fine-grained spatial information across the network, aiding in the precise localisation needed for accurate synthesis. Mean Squared Error (MSE) loss was utilised to train the UNet, which measures the average squared difference between the estimated values and what is estimated.
@@ 3.2.2 GAN Models
GANs consist of a Generator $G$ that generates fake data from random noise $z$ and a Discriminator $D$ that distinguishes real from fake data. They are trained simultaneously in a minimax game where $G$ tries to minimise $\log(1-D(G(z)))$ and $D$ tries to maximise $\log D(x)+\log(1-D(G(z)))$. This adversarial process results in $G$ producing data increasingly similar to the real data. We have implemented two variations of GANs differentiated by how they incorporate the input BF image condition:
@@ 3.2.3 Diffusion Models
Diffusion models simulate the gradual transformation of data from a structured state into a random state and then learn to reverse this process to generate new data samples. The mathematical framework of diffusion models involves two key processes: the forward diffusion process, which adds noise to the data, and the reverse process, which learns to generate data by removing noise.
The forward process is modeled as a Markov chain that starts with the original data $x_{0}$ and adds noise at each step $t$, resulting in a sequence of increasingly noisy data $x_{1},x_{2},\ldots,x_{T}$, where $T$ is the total number of diffusion steps. Mathematically, the transition from $x_{t-1}$ to $x_{t}$ can be defined using a Gaussian distribution:
$x_{t}=\sqrt{1-\beta_{t}}\,x_{t-1}+\sqrt{\beta_{t}}\,\epsilon_{t},$
where $\beta_{t}$ is a variance schedule that determines the amount of noise added at each step, and $\epsilon_{t}\sim\mathcal{N}(0,I)$ is a sample from a standard Gaussian distribution.
The reverse process is defined by a sequence of conditional distributions, each aiming to predict the clean data $x_{t-1}$ given the noisy data $x_{t}$:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}(x_{t},t)),$
where $\mu_{\theta}(x_{t},t)$ and $\sigma_{\theta}(x_{t},t)$ are functions parameterized by the model parameters $\theta$, predicting the mean and variance of the Gaussian distribution for $x_{t-1}$ based on $x_{t}$.
Palette [21]. Within the diffusion model framework, Palette utilises an attention U-Net architecture for the denoising process; the input BF images are concatenated with noisy images, providing conditional context throughout the iterative denoising steps. As with the UNet implementation, MSE loss is used to guide the denoising process towards producing high-quality IF images.
@@ 3.3 Evaluation Methods
Various algorithms have been proposed in previous studies to evaluate synthetic images [11, 22, 2]. In this paper, we define two categories of metrics: image quality evaluation metrics and biological meaning evaluation metrics. Several images near the edges of the plates contained no useful information because the corresponding ground truth IF images for these regions were entirely black. To ensure more accurate analysis, we excluded these non-informative images from the dataset.
@@ 3.3.2 Biological Meaning Metrics: Feature Correlation and MoA Prediction Accuracy
The first evaluation metric is the correlation between these features. The mean values for each feature group were organized into correlation matrices and visualised as heatmaps. In CellProfiler, features from each object were divided into several categories: area/shape, colocalisation, granularity, intensity, neighbours, radial distribution, and texture.
The second evaluation metric is the predictive performance of these single-cell features for determining the mechanism of action (MoA) of perturbations e.g. compounds. The MoA refers to the biological process through which a compound achieves its pharmacological effect, such as targeting specific proteins or modulating certain pathways. Understanding a compound’s MoA is critical in pre-clinical phases as it helps identify potential efficacy and toxicity effects. Synthetic IF images can be used for automatic diagnosis [28], such as identifying diseased cells. However, we did not explore this due to the limited number of cell lines.
@@ 4.1 Pix2Pix Model Performs the Best Regarding Image Quality Evaluation
Visual Comparison. We invited two biologists who work closely with IF staining to visually compare the performance of different models. Both agreed that through sufficient training, all five models captured the distinctive style of the IF images, though, there are noticeable differences in details between the generated images and the ground truth images. Among these models, we discovered that the Pix2Pix model captures the details of the real IF images the most accurately.
@@
In Figure 3(a), when comparing the MSE values across different channels, we noticed that, with the exception of SPADE, the MSEs for the DNA channel in all models are significantly lower ($p<0.05$) than those for other channels. This suggested that predicting the cell nucleus channel is easier to achieve than transferring the styles of BF or searching for subtle and embedded subcellular structures in other channels (Mito, AGP, ER, and RNA). This is likely because nuclei have a clearer shape and more defined semantics compared to other structures. To improve synthesis across all structures, additional conditional mechanisms may be required during the synthesis process across all model families.
@@ 4.3 Acceptable Time and Computational Cost of Synthetic IF Images Compared to Bench Work
In addition to the time cost, here we compared the financial costs related to both wet lab cell painting and virtual cell painting image synthesis. It is worth nothing that since the learning curves of researchers is both difficult and subjective to evaluate, we opt only to consider hardware and consumable – not human – resource costs in our comparisons.
@@
4.4 Generalisability to Unseen Circumstances
4.4.1 Lack of Generalisability to Unseen Cell Lines
4.4.2 Generalisability to Unseen Pertubations
In our experiments, we used microscopy images with compound treatments. However, gene editing can also perturb cells and alter their morphological features across different channels. To evaluate the generalisability of generative AI on unseen perturbations, we used our best-performing model, Pix2Pix, which was trained on BF-IF image pairs with compound treatments, and tested it on gene-edited A549 and U2OS cell lines.
Although we discovered a significant performance drop in synthetic image quality (MSE 37.57, STD 20.93, vs. 29.29, STD 16.60; mean increase 8.28), this increase is still not as significant as the difference between cell lines (mean drop 12.08) shown in Figure 6.
In 2024, Chandrasekaran et al. [4] established a correlation between chemical (compound) and genetic (gene-editing) perturbations based on morphological features extracted from CellProfiler. They identified that compound treatments can cause similar IF image appearances as gene editing, producing comparable morphological features. Subsequently, in this section, we investigated whether a model trained solely on compound-treated BF and IF images could accurately predict the IF images of gene-edited cells based on their BF images. Specifically, we wanted to see if the model could identify similar compound-gene pairs as observed in real images. Thus, we used CellProfiler to extract features from both real and synthetic images, aggregated the median values of each well, and used the t-SNE algorithm to compress the CellProfiler features into two dimensions.
In the study by Chandrasekaran et al., the top positively correlated gene–compound match in A549 cells is the PLK1 inhibitor compound BI-2536, which matched with CRISPR against PLK1, followed by the Aurora kinase inhibitor AMG900, which matched with CRISPR against AURKB. We highlighted these compound-CRISPR pairs in Figure 7. We noticed that these highly similar pairs identified by real image features (red and orange dots in Figure 7 (2)) also exhibit high similarity in features extracted from synthetic images (red and orange dots in Figure 7 (1)). This underscores the generalizability of the Pix2Pix model to unseen perturbations and its ability to summarize biological phenomena as effectively as real images.
Figure 7: T-SNE visualization of CellProfiler features extracted from synthetic A549 IF images generated by Pix2Pix (1) and from real A549 images (2). Triangles represent gene-edited images, and circles represent compound-treated images. Highlighted in both panels are the top positively correlated gene–compound matches identified by [4]. Both AMG900 (chemical) and orantinib (chemical) target AURKB. BI-2536 (chemical) is a PLK1 inhibitor compound, targeting PLK1 (gene).
@@ 5.1 Three Typical Artifacts in Synthetic IF Images
In Figure 2, we highlighted three types of artifacts that were identified when generative AI predicted IF images from BF images: 1) Blurring in UNet generated images (Figure 2c): This artifact is prevalent in the UNet-predicted IF images through all channels. This is possibly attributed to the model’s MSE-based optimisation loss function. The MSE loss encourages minimising pixel-wise differences, blurring fine details and leading to overly smoothed images. Blurring can obscure critical cellular details, making it challenging to identify and analyse fine structures within the cells. 2) Collapsed inner structure especially in the AGP channel: Some models fail to accurately replicate the structure, creating hallucinations or "collapsed biological structures." This hallucination is likely due to the model’s inability to condition the synthesis procedure upon given BF conditions. The presence of hallucinated or collapsed structures can mislead researchers about the actual actin, Golgi, plasma membrane (AGP) structures present in the sample, resulting in incorrect conclusions about cellular organization and function. 3) Intensity shift and high background noises: In diffusion models, the predicted images often have significant background noise and intensity shifts. These artifacts can occur due to the nature of the diffusion process, where each step requires intensity clipping that may introduce noise and modify image brightness. This intensity shift can complicate the quantification of fluorescence intensity, leading to potential errors in measuring protein expression levels, cell viability, or other quantitative metrics.
@@ 5.3 Problematic MSE
where $y_{i}$ is the ground truth pixel value and $\hat{y}_{i}$ is the predicted pixel value. However, MSE loss is limited as an evaluation metric as it only considers the pixel value differences and aggregates the whole image regardless of background and signals. In a biological context, this may pose serious potential issues in downstream applications.
@@
In Figure 9, we present two synthetic IF images (2,3) alongside their corresponding ground truth image (1). Figure 9 (2a) was generated by the Pix2Pix network while Figure 9 (3a) was produced by intensity reduction. As shown in Figure 9 (2b), while Figure 9 (2a) has a smaller image pixel intensity difference compare to the real image, and therefore a lower MSE loss, the cell morphology (in terms of nuclear size, shape and texture) is incorrectly predicted, leading to an inaccurate estimation of cell mortality, shown with two red arrows. On the other hand, Figure 9 (3b) largely retains nuclear morphology despite large differences in stain intensity compared to the real image. Thus, while MSE, PSNR, and SSIM can roughly assess image intensity, they do not fully capture the biological relevance. As such, leveraging metrics that assess the biological significance of these synthetic images is crucial for robust model evaluation.
@@ 5.4 Discrepancy in Feature Correlation and MoA Prediction Accuracy
As illustrated, the single-cell features identified in UNet and GAN-based models closely follow the real feature manifold, which is consistent with the image quality results. In contrast, the features of cells generated by diffusion models occupy a different space and exhibit different characteristics. However, this does not imply that these out-of-real data points lack classification ability. Indeed, we observed a significant location difference between the DMSO group and the inhibitor-treated group in features generated by the diffusion models, as highlighted by the red and purple points in Figure 10. This indicates the potential of synthetic IF in predicting drug response and in high-content drug screening. The low feature correlation may harm the explainability of the classification model, but the synthetic images can still contribute valuable information for distinguishing between different treatment groups.